OpenStudy (anonymous):

If the product of 14 and a number is decreased by 7, and then the result is divided by 5, the answer is 7. What is the unknown number

OpenStudy (johnweldon1993):

Let the number = x \[\frac{ 14x - 7 }{ 5 } = 7\] First...multiply both sides by 5...to get rid of that fraction 14x - 7 = 35 What next...?

OpenStudy (anonymous):

the unknown number would be 3? (:

OpenStudy (johnweldon1993):

That would be correct! :)